A recipe for applesauce calls for 1/3 cup of honey. The recipe makes 8 servings. How many cups of honey are needed to make 20 servings?
Answer by checkley79(3341)

.33/8=x/20 cross multiply.
8x=.33*20
8x=6.67
x=6.67/8
x=.8325 cups of honey is needed.